If X is a binomial random variable with expected value 6 and variance 2.4. Find the
probability of X=5.
1
Expert's answer
2020-04-02T13:47:52-0400

The expected value np=6

Variance npq=2.4

q="\frac{npq} {np} =\frac {2.4}{6}" =0.4

p=1-0.4=0.6

np=6=n*0.6

n="\frac {6}{0.6}" =10

P(X=5)

"P(X=x) =\binom {n} {x} p^x (1-p)^{n-x}"

"P(X=5) =\binom {10}{5}*0.6^5 *0.4^5"

=0.2007